# Towards Efficient Real-Time Video Motion Transfer via Generative Time Series Modeling

### v1

@@ 3.2.3 Variational Recurrent Neural Network (VRNN)
A VRNN is a generative time series model which combines the sequential modeling ability of RNNs with the probabilistic latent structure of VAEs. The VRNN model is essentially an RNN that contains a VAE at each time step [8]. In order to maintain consistency among the various deep learning models in our work, we have used GRU as the RNN unit of the VRNN. The key steps describing how a VRNN works are descibed as follows.
For a time series of keypoints, the VRNN maintains a hidden state $\bm{h}_{t}$ at each timestep $t$, which is updated based on the input $\bm{x}_{t}$, the prior $\bm{z}_{t}$, and the previous hidden state $\bm{h}_{t-1}$. The hidden state update in a VRNN is given as below:
$\bm{h}_{t}=f\left(\varphi_{\tau}^{\bm{x}}(\bm{x}_{t}),\varphi_{\tau}^{\bm{z}}(\bm{z}_{t}),\bm{h}_{t-1};\theta\right),$ (13)
where $f$ represents a nonlinear activation function, and $\varphi_{\tau}^{\bm{x}}(\cdot)$ and $\varphi_{\tau}^{\bm{z}}(\cdot)$ are neural network functions parameterized by $\tau$ for input and latent variable transformations respectively. In contrast with a VAE, the latent variable $\bm{z}_{t}$ in case of a VRNN is sampled from a posterior distribution conditioned on both the input and the hidden state as below:
$q(\bm{z}_{t}|\bm{x}_{t},\bm{h}_{t-1})=\mathcal{N}\left(\mu_{\bm{z},t},\sigma_{\bm{z},t}^{2}\right),$ (14)
where $[\mu_{\bm{z},t},\sigma_{\bm{z},t}]=\varphi_{\tau}^{\text{enc}}(\varphi_{\tau}^{\bm{x}}(\bm{x}_{t}),\bm{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{enc}}$ is the neural network for encoder, and $\mu_{\bm{z},t}$ and $\sigma_{\bm{z},t}$ are the parameters of the posterior distribution.
Similar to a VAE, the prior distribution in case of a VRNN over $\bm{z}_{t}$ is also modeled as a Gaussian but in this case it depends on the previous hidden state of the GRU as below:
$p(\bm{z}_{t}|\bm{h}_{t-1})=\mathcal{N}\left(\mu_{0,t},\sigma_{0,t}^{2}\right),$ (15)
where $[\mu_{0,t},\sigma_{0,t}]=\varphi_{\tau}^{\text{prior}}(\bm{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{prior}}$ is the neural network for prior distribution, $\mu_{\bm{0},t}$ and $\sigma_{\bm{0},t}$ are the parameters of the prior distribution. The generative process reconstructs $\bm{x}_{t}$ from $\bm{z}_{t}$ and $\bm{h}_{t-1}$:
$p(\bm{x}_{t}|\bm{z}_{t},\bm{h}_{t-1})=\mathcal{N}\left(\mu_{\bm{x},t},\sigma_{\bm{x},t}^{2}\right),$ (16)
where $[\mu_{\bm{x},t},\sigma_{\bm{x},t}]=\varphi_{\tau}^{\text{dec}}(\varphi_{\tau}^{\bm{z}}(\bm{z}_{t}),\bm{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{dec}}$ is the neural network for decoder, and $\mu_{\bm{x},t}$ and $\sigma_{\bm{x},t}$ are the parameters of the decoder output distribution.
@@
$\mathcal{L}_{\text{VRNN}}=\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}\mathbb{E}_{q(\bm{z}_{\leq T}|\bm{x}_{\leq T})}\Bigg{[}\sum_{t=1}^{T}\Big{(}-KL\big{(}q(\bm{z}^{(i)}_{t}|\bm{x}^{(i)}_{\leq t},\bm{z}^{(i)}_{<t})$ (17)
$\|\,p(\bm{z}^{(i)}_{t}|\bm{x}^{(i)}_{<t},\bm{z}^{(i)}_{<t})\big{)}+\log p(\bm{x}^{(i)}_{t}\mid\bm{z}^{(i)}_{\leq t},\bm{x}^{(i)}_{<t})\Big{)}\Bigg{]}\Bigg{]}$
@@
During inference, similar to a GRU, the VRNN takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and recursively generates the next $k$ future values using its predictions from previous timesteps as inputs in an autoregressive manner. Figure 4 provides a graphical representation of all the computations for a VRNN [55] .
@@ 3.2.4 Normalizing Flow (NF)
Normalizing Flows (NF) are a class of generative models that transform a simple base distribution (e.g., Gaussian) into a complex target distribution using a sequence of differentiable and invertible mappings, enabling both efficient sampling and exact likelihood evaluation [38]. For the invertible mapping, NFs apply the ‘change of variables’ rule [45] which is described as follows. Let $\bm{x}\in\bm{X}$ be the observed data with a probability distribution $p_{\bm{x}}$ which in our case is a time series of keypoints. Let the corresponding latent variable be $\bm{z}\in\bm{Z}$ with a probability distribution $p_{\bm{z}}$, and let g be an invertible and differentiable function where $\bm{X}=g(\bm{Z})$ and f = $g^{-1}$. Given the probability density function of $\bm{z}$ , the change of variables formula for calculating the probability density function of $\bm{x}$ is as follows:
$p_{X}(\bm{x})=p_{Z}(f(\bm{x}))\left|\det\left(\frac{\partial f(\bm{x})}{\partial\bm{x}^{T}}\right)\right|$ (18)
$\log(p_{X}(\bm{x}))=\log\left(p_{Z}(f(\bm{x}))\right)+\log\left(\left|\det\left(\frac{\partial f(\bm{x})}{\partial\bm{x}^{T}}\right)\right|\right)$ (19)
where $p_{Z}=\mathcal{N}(z;\mu=z,\sigma^{2}=1)$ and $\frac{\partial f(\bm{x})}{\partial\bm{x}^{T}}$ is the Jacobian of f at $\bm{x}$.
For practical applications involving high dimensional data, estimating the Jacobian becomes computationally infeasible. To address these limitations, deep neural network (DNN) based estimation has been proposed for efficient and tractable computation of Jacobians. Real-valued Non-Volume Preserving (RealNVP) transformations, which use coupling layers as proposed by [11], are a type of DNN-based approach that we use in our work to perform normalizing flow transformations. The equations of coupling layers used in Real NVP are described as follows.
Let $\bm{x}$ be a D dimensional input data to the NF where $d<D$, and $\bm{y}$ be the output of a coupling layer. The transformation performed by the coupling layer is defined by the following equations:
$\bm{y}_{1:d}=\bm{x}_{1:d}$ (20)
$\bm{y}_{d+1:D}=\bm{x}_{d+1:D}\odot\exp\left(s(\bm{x}_{1:d})\right)+t(\bm{x}_{1:d})$ (21)
where $\odot$ is the hadamard (element-wise) product, and s() and t() are scale and translation functions that are feedforward neural networks respectively from $\mathbb{R}^{d}\rightarrow\mathbb{R}^{D-d}$. The Jacobian matrix becomes $\exp\left(\sum_{j}s\left(\bm{x}_{1:d}\right)_{j}\right)$.
@@
$\bm{x}_{1:d}=\bm{y}_{1:d}$ (22)
$\bm{x}_{d+1:D}=\left(\bm{y}_{d+1:D}-t(\bm{y}_{1:d})\right)\odot\exp\left(-s(\bm{y}_{1:d})\right)$ (23)
[…]
@@
For splitting the input as shown in Eqs. 20 and 21, channel-wise masking is applied in a shifting pattern [11]. A number of coupling layers are applied to map $\bm{x}\rightarrow\bm{y}_{1}\rightarrow\cdots\rightarrow\bm{y}_{k-1}\rightarrow\bm{z}$ all the while alternating the dimensions. The part that is identical in the current coupling layer is updated in the next one. Figure 5 shows the shifting pattern of applying the coupling layer equations to the two parts of the input.
In our work, similar to the approach followed for the VAE, given $k>1$ as the prediction horizon, the NF is trained to predict ${\bm{x}_{t+k}}$ from ${\bm{x}_{t}}$ at time $t$ over $N$ training samples using the loss function as given below:
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\bm{x}}_{t+k}^{(i)}-\bm{x}_{t+k}^{(i)})^{2}$ (25)
$\mathcal{L}_{\text{NLL}}=-\frac{1}{N}\sum_{i=1}^{N}\log p_{\bm{x}}(\bm{x}^{(i)}_{t})$ (26)
The first term in the loss function minimizes the reconstruction loss over $N$ training samples between the predicted output $\bm{\hat{x}}$ and its corresponding ground truth $\bm{x}$ at time $t+k$ while the second term is the estimated NLL of the data.
During inference similar to a VAE, the NF takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and generates the next $k$ future values by reconstructing each ${\bm{x}_{k+i}}$ from ${\bm{x}_{i}}$, for $i=1,\dots,n$, where $k>=n$, thereby shifting the input index forward with each prediction.
@@ 3.2.5 Gated Recurrent Unit-Normalizing Flow (GRU-NF)
GRU-NF is a generative time series model proposed in [42] which combines the sequential modeling ability of GRUs with the exact likelihood modeling capability of NFs. The GRU-NF aims to effectively capture both the sequential structure and the distributional complexity of the underlying data. In our implementation, we use RealNVP as the NF component in the GRU-NF architecture. The key steps stating the details of how a GRU-NF works is described as follows.
For a time series of D dimensional keypoints ${\bf x}$, the GRU takes $\bm{x}_{1:T}$ and generates a sequence of hidden states over this time window of $1:T$ as below:
$\{\bm{h}_{1},\ldots,\bm{h}_{T}\}=\text{GRU}(\{\bm{x}_{1},\ldots,\bm{x}_{T}\};\bm{h}_{0})$ (27)
Here the cold-start hidden state $\bm{h}_{{0}}$ is set to $\vec{0}$. Note that since the hidden state $\bm{h}$ contains the temporal dependencies from previous timesteps, it serves as a conditioning variable for the NF model to generate samples in its latent space as below.
$\bm{z}_{1:T}=f(\bm{x}_{1:T}|\bm{h}_{1:T})$ (28)
@@
$\log p_{\bm{x}}(\bm{x}_{1:T}|\bm{h}_{1:T};\theta)=\log p_{\bm{z}}(f(\bm{x}_{1:T}|\bm{h}_{1:T}))$
$+\log\left|\det J_{f}(\bm{x}_{1:T}|\bm{h}_{1:T})\right|$ (29)
@@
To calculate $\log p_{\bm{x}}(\bm{x}|\mathbf{h};\theta)$, the hidden state information from the GRU is concatenated to the inputs of the scaling and translation functions in the coupling layers of the NF i.e. $s(\mathrm{concat}(\bm{x}_{1:d},\bm{h}))$ and $t(\mathrm{concat}(\bm{x}_{1:d},\bm{h}))$, where $s(\cdot)$ and $t(\cdot)$ represent the scaling and translation functions, respectively, and $\bm{x}_{1:d}$ denotes part of the input vector where $d<D$.
@@
$\mathcal{L}_{GRU-NF}=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\log p_{\bm{x}}(\bm{x}_{t}^{(i)}|\bm{h}_{t}^{(i)};\theta)$ (30)
During inference, the GRU is applied on the past input $\bm{x}_{t-1}$ and hidden state $\bm{h}_{t-1}$ from the previous time step as follows:
$\bm{h}_{t}=\text{GRU}(\bm{x}_{t-1};\bm{h}_{t-1})$ (31)
Then, by sampling a noise vector $\bm{z}_{t}\in\mathbb{R}^{D}$ from an isotropic Gaussian distribution, we apply the inverse flow to obtain a sample of the time series at the next time step:
${\bm{x}}_{t}=f^{-1}(\bm{z}_{t}\mid{\bm{h}}_{t})$ (32)
We then use this newly predicted output to compute the next hidden state ${\bm{h}}_{t+1}$ via the GRU, and repeat this process over the desired inference horizon. The flow diagram of the components of GRU-NF architecture is displayed in Figures 6 and 7.
$h_{0}$$x_{1:T}$$h_{1:T}$$z_{1:T}$(27)(27)(28)(28)

### v2

@@ 3.2.3 Variational Recurrent Neural Network (VRNN)
A VRNN is a generative time series model which combines the sequential modeling ability of RNNs with the probabilistic latent structure of VAEs. The VRNN model is essentially an RNN that contains a VAE at each time step [Chung2015]. In order to maintain consistency among the various deep learning models in our work, we have used GRU as the RNN unit of the VRNN. The key steps describing how a VRNN works are descibed as follows.
For a time series of keypoints, the VRNN maintains a hidden state $\boldsymbol{h}_{t}$ at each timestep $t$, which is updated based on the input $\boldsymbol{x}_{t}$, the prior $\boldsymbol{z}_{t}$, and the previous hidden state $\boldsymbol{h}_{t-1}$. The hidden state update in a VRNN is given as below:
$\boldsymbol{h}_{t}=f\left(\varphi_{\tau}^{\boldsymbol{x}}(\boldsymbol{x}_{t}),\varphi_{\tau}^{\boldsymbol{z}}(\boldsymbol{z}_{t}),\boldsymbol{h}_{t-1};\theta\right),$ (13)
where $f$ represents a nonlinear activation function, and $\varphi_{\tau}^{\boldsymbol{x}}(\cdot)$ and $\varphi_{\tau}^{\boldsymbol{z}}(\cdot)$ are neural network functions parameterized by $\tau$ for input and latent variable transformations respectively. In contrast with a VAE, the latent variable $\boldsymbol{z}_{t}$ in case of a VRNN is sampled from a posterior distribution conditioned on both the input and the hidden state as below:
$q(\boldsymbol{z}_{t}|\boldsymbol{x}_{t},\boldsymbol{h}_{t-1})=\mathcal{N}\left(\mu_{\boldsymbol{z},t},\sigma_{\boldsymbol{z},t}^{2}\right),$ (14)
where $\mu_{\boldsymbol{z},t},\sigma_{\boldsymbol{z},t}]=\varphi_{\tau}^{\text{enc}}(\varphi_{\tau}^{\boldsymbol{x}}(\boldsymbol{x}_{t}),\boldsymbol{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{enc}}$ is the neural network for encoder, and $\mu_{\boldsymbol{z},t}$ and $\sigma_{\boldsymbol{z},t}$ are the parameters of the posterior distribution.
Similar to a VAE, the prior distribution in case of a VRNN over $\boldsymbol{z}_{t}$ is also modeled as a Gaussian but in this case it depends on the previous hidden state of the GRU as below:
$p(\boldsymbol{z}_{t}|\boldsymbol{h}_{t-1})=\mathcal{N}\left(\mu_{0,t},\sigma_{0,t}^{2}\right),$ (15)
where $\mu_{0,t},\sigma_{0,t}]=\varphi_{\tau}^{\text{prior}}(\boldsymbol{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{prior}}$ is the neural network for prior distribution, $\mu_{\boldsymbol{0},t}$ and $\sigma_{\boldsymbol{0},t}$ are the parameters of the prior distribution. The generative process reconstructs $\boldsymbol{x}_{t}$ from $\boldsymbol{z}_{t}$ and $\boldsymbol{h}_{t-1}$:
$p(\boldsymbol{x}_{t}|\boldsymbol{z}_{t},\boldsymbol{h}_{t-1})=\mathcal{N}\left(\mu_{\boldsymbol{x},t},\sigma_{\boldsymbol{x},t}^{2}\right),$ (16)
where $\mu_{\boldsymbol{x},t},\sigma_{\boldsymbol{x},t}]=\varphi_{\tau}^{\text{dec}}(\varphi_{\tau}^{\boldsymbol{z}}(\boldsymbol{z}_{t}),\boldsymbol{h}_{t-1})$. Here, $\varphi_{\tau}^{\text{dec}}$ is the neural network for decoder, and $\mu_{\boldsymbol{x},t}$ and $\sigma_{\boldsymbol{x},t}$ are the parameters of the decoder output distribution.
@@
$\mathcal{L}_{\text{VRNN}}=\frac{1}{N}\sum_{i=1}^{N}\Bigg[\mathbb{E}_{q(\boldsymbol{z}_{\leq T}|\boldsymbol{x}_{\leq T})}\Bigg[\sum_{t=1}^{T}\Big(-KL\big(q(\boldsymbol{z}^{(i)}_{t}|\boldsymbol{x}^{(i)}_{\leq t},\boldsymbol{z}^{(i)}_{<t})$ (17)
$\|\,p(\boldsymbol{z}^{(i)}_{t}|\boldsymbol{x}^{(i)}_{<t},\boldsymbol{z}^{(i)}_{<t})\big)+\log p(\boldsymbol{x}^{(i)}_{t}\mid\boldsymbol{z}^{(i)}_{\leq t},\boldsymbol{x}^{(i)}_{<t})\Big)\Bigg$$\Bigg]$
@@
During inference, similar to a GRU, the VRNN takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and recursively generates the next $k$ future values using its predictions from previous timesteps as inputs in an autoregressive manner. Figure 4 provides a graphical representation of all the computations for a VRNN [Ullah2020] .
@@ 3.2.4 Normalizing Flow (NF)
Normalizing Flows (NF) are a class of generative models that transform a simple base distribution (e.g., Gaussian) into a complex target distribution using a sequence of differentiable and invertible mappings, enabling both efficient sampling and exact likelihood evaluation [Papamakarios2021]. For the invertible mapping, NFs apply the ‘change of variables’ rule [Rezende2015] which is described as follows. Let $\boldsymbol{x}\in\boldsymbol{X}$ be the observed data with a probability distribution $p_{\boldsymbol{x}}$ which in our case is a time series of keypoints. Let the corresponding latent variable be $\boldsymbol{z}\in\boldsymbol{Z}$ with a probability distribution $p_{\boldsymbol{z}}$, and let g be an invertible and differentiable function where $\boldsymbol{X}=g(\boldsymbol{Z})$ and f = $g^{-1}$. Given the probability density function of $\boldsymbol{z}$ , the change of variables formula for calculating the probability density function of $\boldsymbol{x}$ is as follows:
$p_{X}(\boldsymbol{x})=p_{Z}(f(\boldsymbol{x}))\left|\det\left(\frac{\partial f(\boldsymbol{x})}{\partial\boldsymbol{x}^{T}}\right)\right|$ (18)
$\log(p_{X}(\boldsymbol{x}))=\log\left(p_{Z}(f(\boldsymbol{x}))\right)+\log\left(\left|\det\left(\frac{\partial f(\boldsymbol{x})}{\partial\boldsymbol{x}^{T}}\right)\right|\right)$ (19)
where $p_{Z}=\mathcal{N}(z;\mu=z,\sigma^{2}=1)$ and $\frac{\partial f(\boldsymbol{x})}{\partial\boldsymbol{x}^{T}}$ is the Jacobian of f at $\boldsymbol{x}$.
For practical applications involving high dimensional data, estimating the Jacobian becomes computationally infeasible. To address these limitations, deep neural network (DNN) based estimation has been proposed for efficient and tractable computation of Jacobians. Real-valued Non-Volume Preserving (RealNVP) transformations, which use coupling layers as proposed by [Dinh2016], are a type of DNN-based approach that we use in our work to perform normalizing flow transformations. The equations of coupling layers used in RealNVP are described as follows.
Let $\boldsymbol{x}$ be a D dimensional input data to the NF where $d<D$, and $\boldsymbol{y}$ be the output of a coupling layer. The transformation performed by the coupling layer is defined by the following equations:
$\boldsymbol{y}_{1:d}=\boldsymbol{x}_{1:d}$ (20)
$\boldsymbol{y}_{d+1:D}=\boldsymbol{x}_{d+1:D}\odot\exp\left(s(\boldsymbol{x}_{1:d})\right)+t(\boldsymbol{x}_{1:d})$ (21)
where $\odot$ is the hadamard (element-wise) product, and s() and t() are scale and translation functions that are feedforward neural networks respectively from $\mathbb{R}^{d}\rightarrow\mathbb{R}^{D-d}$. The Jacobian matrix becomes $\exp\left(\sum_{j}s\left(\boldsymbol{x}_{1:d}\right)_{j}\right)$.
@@
$\boldsymbol{x}_{1:d}=\boldsymbol{y}_{1:d}$ (22)
$\boldsymbol{x}_{d+1:D}=\left(\boldsymbol{y}_{d+1:D}-t(\boldsymbol{y}_{1:d})\right)\odot\exp\left(-s(\boldsymbol{y}_{1:d})\right)$ (23)
$x_{1:d}\odot$$x_{d+1:D}=t,s\odot$
@@
For splitting the input as shown in Eqs. 20 and 21, channel-wise masking is applied in a shifting pattern [Dinh2016]. A number of coupling layers are applied to map $\boldsymbol{x}\rightarrow\boldsymbol{y}_{1}\rightarrow\cdots\rightarrow\boldsymbol{y}_{k-1}\rightarrow\boldsymbol{z}$ all the while alternating the dimensions. The part that is identical in the current coupling layer is updated in the next one. Figure 5 shows the shifting pattern of applying the coupling layer equations to the two parts of the input.
In our work, similar to the approach followed for the VAE, given $k>1$ as the prediction horizon, the NF is trained to predict $\boldsymbol{x}_{t+k}$} from $\boldsymbol{x}_{t}$} at time $t$ over $N$ training samples using the loss function as given below:
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\boldsymbol{x}}_{t+k}^{(i)}-\boldsymbol{x}_{t+k}^{(i)})^{2}$ (25)
$\mathcal{L}_{\text{NLL}}=-\frac{1}{N}\sum_{i=1}^{N}\log p_{\boldsymbol{x}}(\boldsymbol{x}^{(i)}_{t})$ (26)
The first term in the loss function minimizes the reconstruction loss over $N$ training samples between the predicted output $\boldsymbol{\hat{x}}$ and its corresponding ground truth $\boldsymbol{x}$ at time $t+k$ while the second term is the estimated NLL of the data.
During inference similar to a VAE, the NF takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and generates the next $k$ future values by reconstructing each $\boldsymbol{x}_{k+i}$} from $\boldsymbol{x}_{i}$}, for $i=1,\dots,n$, where $k>=n$, thereby shifting the input index forward with each prediction.
@@ 3.2.5 Gated Recurrent Unit-Normalizing Flow (GRU-NF)
GRU-NF is a generative time series model proposed in [Rasul2020] which combines the sequential modeling ability of GRUs with the exact likelihood modeling capability of NFs. The GRU-NF aims to effectively capture both the sequential structure and the distributional complexity of the underlying data. In our implementation, we use RealNVP as the NF component in the GRU-NF architecture. The key steps stating the details of how a GRU-NF works is described as follows.
For a time series of D dimensional keypoints $\bf x$}, the GRU takes $\boldsymbol{x}_{1:T}$ and generates a sequence of hidden states over this time window of $1:T$ as below:
$\{\boldsymbol{h}_{1},\ldots,\boldsymbol{h}_{T}\}=\text{GRU}(\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{T}\};\boldsymbol{h}_{0})$ (27)
Here the cold-start hidden state $\boldsymbol{h}_{{0}}$ is set to $\vec{0}$. Note that since the hidden state $\boldsymbol{h}$ contains the temporal dependencies from previous timesteps, it serves as a conditioning variable for the NF model to generate samples in its latent space as below.
$\boldsymbol{z}_{1:T}=f(\boldsymbol{x}_{1:T}|\boldsymbol{h}_{1:T})$ (28)
@@
$\log p_{\boldsymbol{x}}(\boldsymbol{x}_{1:T}|\boldsymbol{h}_{1:T};\theta)=\log p_{\boldsymbol{z}}(f(\boldsymbol{x}_{1:T}|\boldsymbol{h}_{1:T}))+\log\left|\det J_{f}(\boldsymbol{x}_{1:T}|\boldsymbol{h}_{1:T})\right|$ (29)
@@
To calculate $\log p_{\boldsymbol{x}}(\boldsymbol{x}|\mathbf{h};\theta)$, the hidden state information from the GRU is concatenated to the inputs of the scaling and translation functions in the coupling layers of the NF i.e. $s(\mathrm{concat}(\boldsymbol{x}_{1:d},\boldsymbol{h}))$ and $t(\mathrm{concat}(\boldsymbol{x}_{1:d},\boldsymbol{h}))$, where $s(\cdot)$ and $t(\cdot)$ represent the scaling and translation functions, respectively, and $\boldsymbol{x}_{1:d}$ denotes part of the input vector where $d<D$.
@@
$\mathcal{L}_{GRU-NF}=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\log p_{\boldsymbol{x}}(\boldsymbol{x}_{t}^{(i)}|\boldsymbol{h}_{t}^{(i)};\theta)$ (30)
During inference, the GRU is applied on the past input $\boldsymbol{x}_{t-1}$ and hidden state $\boldsymbol{h}_{t-1}$ from the previous time step as follows:
$\boldsymbol{h}_{t}=\text{GRU}(\boldsymbol{y}_{t-1};\boldsymbol{h}_{t-1})$ (31)
Then, by sampling a noise vector $\boldsymbol{z}_{t}\in\mathbb{R}^{D}$ from an isotropic Gaussian distribution, we apply the inverse flow to obtain a sample of the time series at the next time step:
${\boldsymbol{y}}_{t(NF)}=f^{-1}(\boldsymbol{z}_{t}\mid{\boldsymbol{h}}_{t})$ (32)
We then use this newly predicted output to compute the next hidden state ${\boldsymbol{h}}_{t+1}$ via the GRU, and repeat this process over the desired inference horizon. The flow diagram of the components of GRU-NF architecture is displayed in Figures 6 and 7.
$h_{0}$$x_{1:T}$$h_{1:T}$$z_{1:T}$